what is the slope of 1,11,500 and 4,13,600 in book sales
Answer by nerdybill(7384)

Consider: "1,11500 and 4,13600"
as two points:
(1,11500) and (4,13600)
.
Slope between any two points is given as:
slope = m = rise/run = (y2-y1)/(x2-x1)
slope = (y2-y1)/(x2-x1)
slope = (13600-11500)/(4-1)
slope = (2100)/(3)
slope = 700